### v5

# Nearly-zero large-angle anisotropy of the cosmic microwave background

Abstract
The global isotropy of the universe is analyzed on the scale of the cosmic horizon, using the angular correlation function $C(\Theta)$ of cosmic microwave background (CMB) temperature at large angular separation $\Theta$. Even-parity correlation $C_{\rm even}(\Theta)$ is introduced to obtain a direct, precise measure of horizon-scale curvature anisotropy independent of the unknown dipole, with uncertainty dominated by models of Galactic emission. In maps from Planck, $C_{\rm even}(\Theta)$ at $\Theta\simeq 90^{\circ}\pm 15^{\circ}$ is found to be much closer to zero than previously documented measurements of correlation, with an absolute value three to four orders of magnitude smaller than expected in standard theory. Such a small variation from zero is estimated to occur by chance in a fraction $\simeq 10^{-4.3}$ to $\simeq 10^{-2.8}$ of standard realizations. We consider an alternative interpretation of this result, as a signature of a scale-invariant angular symmetry of cosmological initial conditions. An example of such a symmetry is formulated geometrically, by assuming that quantum fluctuations during inflation have spacelike coherence bounded by compact causal diamonds, and become classical perturbations when world lines cross inflationary horizons. It is argued that this interpretation is inconsistent with the standard theory of initial conditions, but is broadly consistent with other cosmological measurements, and is subject to further tests.
keywords:
cosmology: observations – cosmic background radiation – early Universe – inflation – large-scale structure of Universe – cosmology: theory
pubyear: 2025pagerange: Nearly-zero large-angle anisotropy of the cosmic microwave backgroundLABEL:lastpage
1 Introduction
The angular distribution of the cosmic microwave background (CMB) provides our most precise measurement of the large-scale structure of space-time. The last scattering surface of the CMB lies at a comoving distance close to our causal horizon, and the pattern of temperature on the sky on large angular scales is thought to represent a direct, intact relic of the large-scale distribution of space-time curvature in the initial conditions. On angular scales larger than a few degrees, the dynamics of the system and the propagation of light are determined only by gravity (Sachs & Wolfe, 1967; Bardeen, 1980; Hu & Dodelson, 2002). The CMB on such large angular scales is well known to display a precise symmetry: on average, it is almost exactly isotropic, characterized by the fact that at large angular separations $\Theta$, its angular correlation function $C(\Theta)$, defined in Eqs. (4) and (5) below, has a dimensionless value much smaller than unity.
Both the near-perfect isotropy and the smaller-scale departures from uniformity that lead to large scale cosmic structure can in principle be generated by a causal physical process if there is an early inflationary acceleration of the cosmological scale factor, which allows any two locations to have causal contact at a sufficiently early time. In standard inflation theory, departures from uniformity arise from quantum fluctuations. The standard quantum model of these fluctuations (Baumann, 2011; Weinberg, 2008), based on local effective quantum field theory (QFT), agrees with measurements of cosmic structure over a wide range of scales, including CMB correlations on scales less than a few degrees.
However, this theory does not agree very well with the precise large-scale uniformity measured on larger angular scales. Indeed, it has been realized since the earliest measurements of CMB anisotropy, with the COBE satellite (Bennett et al., 1994; Hinshaw et al., 1996), that the universe is actually much smoother at large $\Theta$ than inflation theory typically predicts. The unexpectedly small magnitude of $C(\Theta)$ at large angular separation was confirmed with higher precision in subsequent studies with WMAP and Planck (Bennett et al., 2003, 2011; Planck Collaboration, 2016, 2020b).
In standard cosmology, the small correlation is attributed to a statistical fluke of our particular realized sky; that is, only a small fraction of realizations that agree with structure on smaller scales are as smooth as the real sky on the largest scales (Copi et al., 2009; Muir et al., 2018). Even though large-angle anisotropy provides the most direct probe of initial conditions, $C(\Theta)$ at $\Theta$ larger than a few degrees, or equivalently the angular power spectrum $C_{\ell}$ at angular wavenumber $\ell$ less than about thirty, are generally disregarded in tests of cosmological models. This practice makes sense within the standard framework of inflationary theory, which predicts many possible realizations of the CMB sky that differ significantly from each other on large angular scales.
In spite of a theoretical expectation that casts doubt on its significance, the structure of CMB anisotropy on the largest scales nevertheless remains a unique phenomenon, which preserves a precisely measurable pattern of cosmic initial conditions from the earliest time and largest distance accessible to us. With this special status as the primary motivation, the current paper aims to measure correlations on large scales, using all-sky CMB temperature maps, up to current limits of systematic measurement uncertainty.
At present, these limits are mainly determined by the best comprehensive all-sky models of Galactic foregrounds, developed by the Planck team. Previous studies have not reached this measurement limit, due to another fundamental systematic uncertainty: because CMB maps have unknown dipole ($\ell=1$) harmonic modes subtracted, their measured $C(\Theta)$ differs from that of the CMB horizon itself (Peebles, 2022).
The current work was partly motivated by an earlier study (Hagimoto et al., 2020), which found that the total $C(\Theta)$ at just one angle, exactly $\Theta=90^{\circ}$, where contributions from the unknown dipole identically vanish, lies in a range remarkably close to zero:
$-0.22\mu{\rm K}^{2}<C(\Theta=90^{\circ})<+2.16\mu{\rm K}^{2}.$ (1)
That measurement showed that the CMB at exactly 90 degrees is hundreds of times smaller than the value in typical standard realizations, and closer to zero than all but $0.52\%$ of them.
In this paper, we study the even-parity part of the angular correlation function $C_{\rm even}(\Theta)$, which gives a direct estimate of true horizon-scale correlations, independent of the unknown dipole or any other model parameter, over a wider range of angles. Our new measurements of $C_{\rm even}(\Theta)$ are plotted for several maps in Fig. (1), together with 100 random realizations of the standard model. From the plot it can be immediately seen that the absolute value of $C_{\rm even}(\Theta)$ over a range of angles near $\Theta\simeq 90^{\circ}$ is much smaller than expected, and differs from zero no more than the different maps differ from each other. Quantitatively, the surprising new result is that the variance of $C_{\rm even}(\Theta)$ from zero in the angular range $\Theta\simeq 90^{\circ}\pm 15^{\circ}$ is three to four orders of magnitude smaller than expected in the standard cosmological model.
The standard interpretation of this new fact, as before, is that our particular horizon just represents a very unlikely statistical fluke, and its small correlations on such large scales are of no physical significance. Such small correlation over a range of $\Theta$ is however hard to dismiss lightly as a statistical anomaly: our rank comparison shows that deviations from zero as small as those in the Planck maps occurs in standard realizations with probabilities that range from $\simeq 10^{-4.3}$ to $\simeq 10^{-2.8}$, depending on the Galactic foreground model.
Over the range $\Theta\simeq 90^{\circ}\pm 15^{\circ}$, the measured isotropy is not just surprising in the context of standard theory, it is extraordinary in absolute terms: to illustrate with one example from the estimates below, the fractional dimensionless correlation residual of the smoothest Planck map (NILC) is $\int C_{\rm even}^{2}/(2.7{\rm K})^{4}\sim 2\times 10^{-26}$. In other physical systems, such very small dimensionless numbers can often be traced to the deep mathematical structure of the system. We are bound to ask whether the measured near-perfect isotropy apparently preserved in the CMB could possibly signify a fundamental symmetry of cosmic initial conditions that is not included in standard theory.
A fundamental symmetry that accounts for vanishing angular correlation would have to be a property of any sky, and any realization of initial conditions. It also needs to account for the specific range of angular scales, $\Theta\simeq 90^{\circ}\pm 15^{\circ}$, where nearly-vanishing even-parity correlation is observed. We show here that it is geometrically possible to formulate such a symmetry, defined by covariant relativistic geometrical relationships, that could account for large-angle $C(\Theta)$ measurements, and at the same time agree with the nearly scale-invariant 3D power spectrum of cosmological perturbations on all scales.
Our formulation is based on a scale-invariant principle of causal coherence widely tested in entangled laboratory quantum systems (Zeilinger, 1999; Vilasini & Renner, 2024): namely, that no correlation can occur between systems contained within completely separate regions of space-time, because they are causally independent. A quantum process that occurs entirely in the future of one event, and entirely in the past of some future event at the same location, can modify physical relationships only within a unique 4D region of space-time bounded by their light cones, called a causal diamond. Quantum processes contained within completely separate causal diamonds do not produce physical correlations with each other. A scale- and conformally-invariant symmetry of angular correlation in cosmic initial conditions could arise from physical potential differences generated on inflationary horizons from quantum fluctuations completely contained within causal diamonds during cosmic inflation.
We show here that in principle, such a causally-coherent process could lead to an angular symmetry on any sky, at any time, similar to that measured in the CMB. It is shown below that geometrically-derived angular boundaries of causal correlations between world lines during inflation coincide with the range of angular separations $\Theta\simeq 90^{\circ}\pm 15^{\circ}$ where we measure the smallest even-parity correlations. The conformal causal relationships that lead to this result do not depend on scale, consistent with a nearly scale-invariant spectrum in 3D. In this range of angular separation, it is possible that causally-coherent primordial angular correlations are not only small in magnitude, they actually vanish, because they are generated independently. The simplicity of this geometrical symmetry makes it possible to test in a model-independent way.
The angular symmetry is formulated here from the standard conformal geometry of classical relativistic cosmology, but it is not compatible with the part of inflationary theory that depends on the standard QFT model for quantum fluctuations and initial perturbations. It should be noted in this regard that the standard formulation of initial conditions in the QFT model has previously been criticized on theoretical grounds (Penrose, 1989; Ellis, 1999; Ijjas & Steinhardt, 2016): for example, an initially perfectly uniform and very large comoving patch of a perfectly uniform 3D background space is simply assumed as an initial condition very early during inflation, and specification of a particular initial vacuum state on this spacelike surface completely determines the final pattern of realized perturbations. Moreover, at a basic physical level, the standard QFT model assumes that field modes evolve coherently over the entire initial patch throughout inflation, even at spacelike separations much larger than the inflationary horizon, in spite of known inconsistencies in QFT on scales larger than causal horizons (Cohen et al., 1999; Hollands & Wald, 2004; Stamp, 2015). If large-angle cosmic correlations preserve unique information about gravitational quantum states on horizons, the exceptional symmetry of CMB isotropy may be more physically profound than it appears to be in the familiar context of QFT-based inflation theory (Hogan, 2019).
The plan of this paper is as follows. In Sec. 2 we describe a new measurement of large-angle even-parity CMB correlations. In Sec. 3, we review the relativistic causal structure of inflationary cosmology, and present a covariant formulation of causal relationships that could define symmetries of angular correlation. In Sec. 4 we compare measurements with standard QFT realizations at angular separations where causal symmetries could lead to zero correlation. In Sec. 5, we summarize the alternative interpretations of the data. An overall summary is presented in Sec. 6. In the Appendix (Sec. 7), we describe in more detail how inflation with causally-coherent symmetry differs from the standard QFT picture, and could modify standard concordance cosmology.
2 Measured large-angle anisotropy
2.1 Angular spectrum and correlation function
In standard notation, the angular pattern of a quantity $Q$ on a sphere, such as scalar potential $\Phi$ or CMB temperature $T$, can be decomposed into spherical harmonics $Y_{\ell m}(\theta,\varphi)$:
$Q(\theta,\varphi)=\sum_{\ell}\sum_{m}Y_{\ell m}(\theta,\varphi)a_{\ell m}.$ (2)
The harmonic coefficients $a_{\ell m}$ then determine the angular power spectrum:
$C_{\ell}=\frac{1}{2\ell+1}\sum_{m=-\ell}^{m=+\ell}|a_{\ell m}|^{2}.$ (3)
The angular correlation function is given by its Legendre transform,
$C(\Theta)=\frac{1}{4\pi}\sum_{\ell}(2\ell+1){C}_{\ell}P_{\ell}(\cos\Theta),$ (4)
where $P_{\ell}$ are Legendre polynomials.
As discussed below, $C(\Theta)$ can be separated into two independent sums with odd and even parity. A sum with only even values of $\ell$ gives the unique even-parity correlation $C_{\rm even}(\Theta)$, which is symmetric around $\Theta=\pi/2$.
The same function can be expressed as an all-sky average
$C(\Theta)=\langle Q_{1}Q_{2}\rangle_{\Theta}$ (5)
for all pairs of points $1,2$ separated by angle $\Theta$, or equivalently, an average over all directions $\vec{\Omega}_{i}$
$C(\Theta)=\langle Q_{\vec{\Omega}_{i}}\bar{Q}_{i\Theta}\rangle_{\vec{\Omega}_{i}}$ (6)
where $\bar{Q}_{i\Theta}$ denotes the average value on a circle of angular radius $\Theta$ with center $\vec{\Omega}_{i}$.
The power spectrum $C_{\ell}$ is the statistical tool generally used for tests of cosmological models. However, $C(\Theta)$ provides a more direct signature of geometrical causal relics of initial conditions. Because it correlates azimuthal averages at a polar angle $\Theta$ (Eq. 6), sharp boundaries of causal correlation within spherical horizon surfaces produce exactly zero $C(\Theta)$ in a range of $\Theta$ determined by the polar angle of their circular intersection. Causal boundaries are not apparent in $C_{\ell}$, even though it contains the same statistical information about the angular distribution.
2.2 Scalar perturbations and large-scale CMB anisotropy
On large angular scales, temperature anisotropy in the CMB is largely determined by primordial scalar curvature perturbations $\Phi$ of the cosmological metric on a thin sphere at the location of the last scattering surface (Sachs & Wolfe, 1967). Apart from the Doppler-induced dipolar anisotropy from local motion, their angular distributions on scales larger than a few degrees are the same:
$\delta T(\theta,\varphi)\propto\Phi(\theta,\varphi),$ (7)
and therefore so are their angular correlations:
$C_{T}(\Theta)\propto C_{\Phi}(\Theta).$ (8)
In this sense, CMB correlation provides a direct measurement of angular symmetries in initial conditions.
Gravity also introduces anisotropy on large angular scales much later during the dark energy dominated era, the late-time integrated Sachs-Wolfe effect (ISW) (Hu & Dodelson, 2002; Francis & Peacock, 2010). This effect is generated by primordial perturbations in 3D, as the CMB light propagates through space on our past light cone. In the linear regime, it is also determined by the invariant local scalar potential $\Phi$ that preserves its original primordial spatial distribution from the end of inflation. However, anisotropy from this effect comes from perturbations at comoving distances much smaller than the last scattering surface, and from correlations of matter at different radii. Thus, the measured total CMB anisotropy depends on angular cross-correlations of $\Phi$ in three comoving spatial dimensions.
In the interpretation below, we will neglect other physical effects, such as radiation transport and Doppler motion at recombination, which do not modify the angular spectrum significantly at spherical harmonics with $\ell\lesssim 30$ (Hu & Dodelson, 2002).
2.3 Data
As explained in Hagimoto et al. (2020), we use all-sky CMB maps made with subtracted models of Galactic emission, in order to minimize correlation artifacts introduced by masks. Our analysis is based on foreground-corrected maps of the CMB temperature based on the fifth and third public release databases of the WMAP and Planck collaborations, respectively. For WMAP, we use the ILC map with and without the fitted monopole included. In the case of Planck, we use several different maps based on different techniques for modeling the Galaxy. Recognizing that the noise properties of the foreground-corrected maps are not well characterized and that the 2-point function is correlated between angles, we use the variation between foreground subtraction methods and experiments as a proxy for correlation function uncertainty. We only compare integrated residuals of measured values and standard model realizations of 2-point correlation functions.
For this paper, we used the python wrapper for the Hierarchical Equal Area isoLatitude Pixelization (HEALPix) scheme (Gorski et al., 2005) on maps at a resolution defined by $N_{\text{side}}=256$. We preprocessed the maps by converting them to this resolution and removing their respective dipole spherical harmonic moments. We conduct all measurements and operations on each map independently.
Figure 1: Correlation functions of sky maps and standard-model realizations. The top panel shows the even parity angular correlation function of CMB temperature. Bold colors show $C_{\rm even}(\Theta)$ of Galaxy-subtracted all-sky maps from the WMAP and Planck satellites, as labeled. For comparison, solid black shows the expectation of the standard model, and fine lines show 100 standard, randomly generated sky realizations. For causally coherent initial conditions, this function should vanish over a symmetric band given by Eq. (19), or $75.52^{\circ}<\Theta<104.48^{\circ}$, which is strikingly approximated by the maps, especially those from Planck data; for these, the variance of correlation is three to four orders of magnitude smaller than that in typical standard realizations, as shown quantitatively in Fig. (4). Bottom panel shows the total correlation with the best-fit dipole restored, assuming a causal shadow that extends over the maximal range tested, $75.52^{\circ}<\Theta<135^{\circ}$ (Eq. 18). For a fair comparison, in this panel each realization has a “mock dipole” correction added to minimize its departure from zero in the shadow region. The maps approximate zero more closely than almost any realization, even with this correction.
2.4 Measured correlation function
The top panel of Fig. (1) shows our main new result: a direct measurement of even-parity CMB angular correlation, which is independent of the unknown dipole. It reveals a simple fact, that the absolute value of $C_{\rm even}(\Theta)$ over a significant range of angles is remarkably close to zero.
In particular, the measured variation of $C_{\rm even}(\Theta)$ from zero in the range $\Theta\simeq 90^{\circ}\pm 15^{\circ}$ is orders of magnitude smaller than previously documented correlations. Its absolute value is comparable with differences between different foreground-subtracted maps, that is, apparently as close to zero as current measurements allow. For the three Planck maps— which arguably provide the most accurate models of Galactic foregrounds— the correlation is nearly indistinguishable from zero on the scale plotted in Fig. (1).
The anisotropy is also strikingly small when compared to the expectations of standard cosmology. To illustrate this comparison, Fig. (1) shows 100 examples of $C_{\rm even}(\Theta)$ produced in standard realizations— that is, anisotropy produced from the same quantum fluctuations that generate cosmic structure on smaller scales.
We now proceed to describe how the observed nearly-zero $C_{\rm even}(\Theta)$ could be interpreted as a causal symmetry of causal initial conditions, and then to quantitative comparisons of this interpretation with the standard picture.
Figure 2: Spacetime diagram illustrating causal relationships betweeen events on world lines $A$ and $B$ spaced a comoving distance $\eta_{0}$ apart. Causal correlations can form where the future light cone of an initial event intersects the past light cone of an observer. In the figure, these relationships are causal diamonds, examples of which are the blue and green diamonds. While the relationships in the figure hold in general for any two world lines, to analyze the situation of the observation of the CMB, we imagine the observer to be on world line $A$ at a time $+\eta_{0}$ and set the time $\eta=0$ to be the end of inflation and shortly after, at the recombination epoch. Thus $\eta_{0}$ is the current horizon distance and the CMB is emitted by a 3-sphere with radius $\eta_{0}$, marked by the red ellipse in the figure and denoted $\mathcal{S}_{A}(\eta_{0})$. The goal is to analyze what events during inflation can, even in principle, generate correlations in the CMB and determine the pattern of the correlations. Because of the causal constraint, not all correlations generated by the standard model for CMB fluctuations are present in this picture. CMB temperature anisotropy on large angular scales is dominated by gravitational perturbations near the last scattering surface, near ${\cal S}_{A}(\eta_{0})$, but also has contributions from distortions arising from later perturbations along the light cone at $\eta>0$, shown as ISW in the figure. Incoming information to the world lines $A$ and $B$ during inflation is bounded by inflationary horizons ${\cal H}_{A}$ and ${\cal H}_{B}$. Shaded 4D regions represent causal diamonds bounded by the comoving 2-spheres ${\cal S}_{A}(\eta_{0}/2)$ (blue) and ${\cal S}_{B}(\eta_{0})$ (green). We posit that coherent quantum fluctuations are bounded by causal diamonds and convert into classical potential differences on inflationary horizons, so a coherent perturbation of ${\cal S}_{A}(\eta_{0}/2)$ forms within the blue causal diamond shown, which starts on $A$ at $-\eta_{0}$; this is the last causal diamond whose fluctuations entangle $A$ with $B$, or with any other point at $|\vec{r}-\vec{r}_{A}|=\eta_{0}$. At locations $\vec{r}$ with radius $|\vec{r}-\vec{r}_{A}|<\eta_{0}/2$, coherent fluctuations generate correlations of scalar potential $\Phi(\vec{r})$ with $\Phi(\vec{r}_{A})$ which are independent of correlations of $\Phi(\vec{r}_{A})$ with $\Phi(\vec{r}_{B})$, so angular correlations vanish in a range of $\Theta$ around $\pi/2$, as shown in Fig. (3).
3 Causal correlations during inflation
3.1 Conformal causal structure
The metric for any homogeneous and isotropic cosmological space-time can be written in conformal coordinates (Baumann, 2011; Weinberg, 2008) as
$ds^{2}=a^{2}(t)[c^{2}d\eta^{2}-d\Sigma^{2}],$ (9)
where $t$ denotes proper cosmic time for any comoving observer, $d\eta\equiv dt/a(t)$ denotes a conformal time interval, and $a(t)$ denotes the cosmic scale factor. For a spatially flat model like that observed, the spatial 3-metric in comoving coordinates is
$d\Sigma^{2}=dr^{2}+r^{2}d\Omega^{2},$ (10)
where $r$ is the comoving radial coordinate, and the angular separation $d\Omega$ in standard polar notation satisfies $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta d\varphi^{2}$. Light cones and causal diamonds are defined by null relationships in comoving conformal coordinates,
$d\Sigma=\pm cd\eta.$ (11)
Thus, in conformal coordinates, cosmological causal relationships throughout and after inflation are the same as those in flat spacetime. We adopt coordinates where $\eta=0$ corresponds to the end of inflationary acceleration. On the large scales studied here, it can also be identified with the CMB last scattering surface. In the following, we set $c=1$. Some key relationships and causal diamonds in this geometry are illustrated in Figure (2).
3.2 Causal bounds on coherence and correlation
Cosmological inflation (Baumann, 2011) was introduced to solve a conceptual problem with initial conditions in classical cosmology, sometimes called the “horizon problem”: as the cosmic expansion slows with time due to normal gravity ($\ddot{a}<0$), causal connections are only possible over smaller comoving regions in the past, so there is no causal mechanism for generating any kind of correlations in the initial conditions.
Inflation solves the main problem by introducing early cosmic acceleration, so that the comoving causal horizon moves closer with time rather than farther away  [^1] An excellent tutorial including visualizations of inflationary conformal space-time, causal horizons, and sky projections can be found at https://www.astro.ucla.edu/~wright/cosmo_04.htm. If the scale factor $a(t)$ undergoes many orders of magnitude of expansion during early acceleration with $\ddot{a}>0$ before some epoch $\eta=0$, even very distant comoving world lines were once in causal contact. This causal relationship is shown in Fig. (2): a comoving world line at any finite radial distance $\eta$ lies within the past light cone or “inflationary horizon” ${\cal H}$ at times earlier than $-\eta$. Fig. (2) shows spatial “footprints” of horizons: comoving spherical surfaces ${\cal S}(\eta)$ that pass through the horizon and out of causal contact at time $-\eta$, and come back into view after inflation at time $+\eta$. All points on a spherical surface ${\cal S}(\eta)$ have a causal connection with an event at its center at $-\eta$, so large-scale homogeneity and isotropy, as assumed in Eq. (9), can in principle be generated by a causal physical process.
Now consider the physical process that generates spatial departures from uniformity caused by quantum fluctuations. In the analysis below, we consider a new causal symmetry that follows from a stronger causal constraint than standard inflationary theory. Suppose that all gravitational quantum fluctuations are causally coherent, in a sense well established from direct measurements of quantum entanglement (Zeilinger, 1999; Vilasini & Renner, 2024): physical correlations are not generated by systems in separate regions of space-time. This constraint requires that physical effects of coherent quantum fluctuation states are spatially compact. Specifically, we posit that quantum curvature fluctuations on any world line interval are entangled nonlocally with other locations only within the compact 4-dimensional region of space-time encompassed by its causal diamond. Put another way, physical correlations are bounded by two-way causal relationships.
We further posit that the inflationary horizon ${\cal H}$ imprints a sharp boundary on coherent quantum fluctuations that create curvature perturbations correlated with that world line: Quantum fluctuations create differences in classical potential from any world line in the future of its inflationary horizon ${\cal H}$. That is, the classical potential difference between world lines forms when they cross each others’ horizons.
As discussed in the Appendix, this hypothesis about how quantum fluctuations convert into classical perturbations differs physically from the freezing of fluctuations in the standard QFT model, where coherent plane waves freeze independently on each comoving scale by synchronous cooling as their wavelengths stretch beyond the horizon scale. In that picture, the final observed classical correlations on ${\cal S}(\eta)$ are fixed by initial data laid down coherently in a region much larger than $\eta$, at a time much earlier than $-\eta$.
This formulation is conformally invariant, so the coherent causal constraint applies to relational perturbations on all comoving length scales. As discussed in the Appendix, there may be observable effects of exotic high-order correlations in the 3D pattern of classical curvature perturbations on smaller scales than the current horizon.
Figure 3: A projection along the time dimension of the relationships in Figure 3. The blue and green disks are the ellipses at $\eta=0$. This projection illustrates the geometrical derivation of the disentanglement angle $\Theta_{D}$ (Eq. 12). The angular radius of the intersection circle on ${\cal S}_{A}(\eta_{0}/2)$, $\pi/2-\Theta_{D}/2$, is the maximum angular separation for correlation with external perturbations in any direction, so the directional average in the correlation function $C(\Theta)$ (Eq. 14) vanishes at $\Theta=90^{\circ}\pm\Theta_{D}/2$ (Eq. 15). At the same time, perturbation differences on ${\cal S}_{B}(\eta_{0})$ are independent of causal restrictions, as in the standard picture, at angular separations $<<\Theta_{D}$. These angular causal relationships are conformally invariant, so they apply to any location, epoch or comoving length scale $\eta_{0}$.
3.3 Scale-invariant angular boundaries of causal correlation
Let us suppose that fluctuations generated in separate causal diamonds are separate quantum systems, and do not correlate with each other. We will now analyze how this 4D relationship maps onto angular correlation, which could lead to finite regions of angular separation in which no correlations occur from entanglement between causal diamonds. We refer to a region of angular separation with no causal entanglement as a “causal shadow.”
It is useful to classify angular boundaries according to geometrical criteria determined by different applications of causal-diamond coherence, which map onto different observational tests. These angular causal relationships are conformally invariant, so they apply on all comoving length scales.
3.3.1 3D causal shadow
As discussed above, suppose that causal correlations are bounded by causal diamonds and inflationary horizons. Fig. (2) shows causal diamonds around two world lines $A$ and $B$, with separation $\eta_{0}$. The causal diamonds on these world lines that begin after $-\eta_{0}$ are disentangled from the other: their fluctuations constitute separate quantum systems, and generate uncorrelated perturbations.
Slices of comoving causal-diamond boundary surfaces in 3D are shown in Fig. (3). The sphere ${\cal S}_{A}(\eta_{0}/2)$ is shown because it represents the smallest horizon footprint around $A$ whose fluctuations are entangled with $B$. Smaller causal diamonds around $A$ intersect ${\cal S}_{B}(\eta_{0})$ at larger angles from $B$, but their perturbations are independent of those at $B$. Thus, the circular intersection of ${\cal S}_{A}(\eta_{0}/2)$ with ${\cal S}_{B}(\eta_{0})$ lies at the largest angular separation from the $B$ direction of locations $\vec{r}$ where perturbations $\Phi(\vec{r})-\Phi(\vec{r}_{A})$ are causally correlated with $\Phi(\vec{r}_{B})-\Phi(\vec{r}_{A})$.
As shown in Fig. (3), the angular radius on ${\cal S}_{B}(\eta_{0})$ of its intersection with ${\cal S}_{A}(\eta_{0}/2)$ is
$\Theta_{D}=2\arcsin(1/4)\simeq 28.96^{\circ},$ (12)
and the angular radius of the same intersection on ${\cal S}_{A}(\eta_{0}/2)$ is
$\pi/2-\Theta_{D}/2.$ (13)
Thus, as viewed from $A$, correlations with perturbations at $B$ vanish at angles separated from the $B$ direction $\vec{\Omega}_{B}$ by more than $\pi/2-\Theta_{D}/2$. That is, values of $\Phi-\Phi_{A}$ at locations closer than $B$ with larger angular separation than Eq. (13) are independent of those at $B$.
The correlation function viewed at $A$ is given by Eq. (6), an all-sky average correlation of $\Phi_{B}$ with $\bar{\Phi}_{B\Theta}$, the average value in a circle of radius $\Theta$ centered on $\vec{\Omega}_{B}$:
$C(\Theta)=\langle\Phi_{B}\bar{\Phi}_{B\Theta}\rangle_{\vec{\Omega}_{B}}.$ (14)
Independence of $\Phi$ at angular separation exceeding $\pi/2-\Theta_{D}/2$ from every direction $\vec{\Omega}_{B}$ then leads to zero angular correlation of correlation over a range of angles symmetric around $\pi/2$:
$C_{\mathrm{3D}}(|\Theta-\pi/2|<\arcsin[1/4])=0.$ (15)
As discussed below, this symmetry applies separately to both the even- and odd- parity components of correlation, which are respectively symmetric and antisymmetric around $\Theta=\pi/2$, as well as their sum. It holds for any value of $\eta_{0}$, so it applies to correlations in the interior of any sphere with points on its boundary.
Importantly for direct null tests of CMB anisotropy, Eq. (14) describes cross-correlation of perturbations $\Phi_{B}$ at $|\vec{r}-\vec{r}_{A}|=\eta_{0}$ with $\bar{\Phi}_{B\Theta}$ at $|\vec{r}-\vec{r}_{A}|<\eta_{0}$. The angular shadow of zero correlation derived from intersection of the two spheres at disentanglement (Eq. 15) applies to foreground perturbations correlated with the last scattering surface, and therefore to gravitational anisotropy generated via the ISW effect. The null symmetry applies to observable CMB correlation independent of the location or epoch of observation: even-parity correlation always vanishes over a range of angles $75.52^{\circ}<\Theta<104.48^{\circ}$.
The effect of causal coherence becomes smaller on small angular scales. At angular separations from $A$ on ${\cal S}_{B}(\eta_{0})$ smaller than $\Theta_{D}$, where angular spectra are generally used for cosmological tests, differences from $\Phi_{A}$ on ${\cal S}_{B}(\eta_{0})$ are independent of $\Phi_{A}-\Phi_{B}$, so perturbations approximate the causally-unconstrained independent phases assumed in the standard inflationary picture. As a result, causal constraints have a negligible effect on angular correlations at small scales, or on the 3D power spectrum for scales small compared to any averaging volume.
3.3.2 Thin-sphere causal shadow
Now consider angular correlations of points constrained to lie at the same comoving distance from $A$. At the time $-\eta_{0}$ when the $AB$ relationship is freezing out— when $B$ passes through ${\cal H}_{A}$— locations on ${\cal S}_{A}(\eta_{0})$ with angular separation from $B$ greater than $\Theta=\pi/3$ lie outside of ${\cal H}_{B}$. These points lie outside the causal diamond surface ${\cal S}_{B}(\eta_{0})$, so their angular causal correlation vanishes:
$C_{\rm tscs}(\Theta>\pi/3)=0.$ (16)
This correlation shadow does not have the same robust application to a precise CMB test as the narrower 3D shadow (Eq. 15): it corresponds to zero angular correlation of potentials only on a thin sphere, so it does not include all gravitational anisotropy from interior perturbations. However, it could define a causal bound on correlations indirectly generated with the antipodal direction, as discussed below.
3.3.3 Correlation at larger angular separation
As shown above, all points on a horizon footprint ${\cal S}_{A}$ have a causal connection with their center $A$. The measured $C(\Theta)$ represents an all-sky average (Eq. 14), so it is modified at all angles by perturbations of $A$ relative to ${\cal S}_{A}$. Subtraction of unobservable monopole and dipole components could indirectly generate apparent angular correlations outside the symmetric interval of the 3D causal shadow (Eq. 15).
We have not derived a model-independent causal symmetry for angular separations outside the 3D causal shadow. Even so, we should explore constraints on correlations that include both odd and even parity components, and that also account for the unobserved dipole, to test whether data is consistent with zero total true correlation over a larger range of angular separations. Since we aim to avoid model-dependent assumptions, the simplest approach is to simply exclude large angular separations. For one test, we will exclude separations within the thin-sphere causal bound for correlations with the antipodal point (Eq. 16), $\Theta>2\pi/3$:
$C_{\rm minimal}([\pi/2-\arcsin(1/4)]<\Theta<2\pi/3)=0.$ (17)
We will also test the possibility of a wider causal shadow,
$C_{\rm maximal}([\pi/2-\arcsin(1/4)]<\Theta<3\pi/4)=0.$ (18)
4 Tests of CMB Symmetries
4.1 Dipole subtraction and parity separation
It is not possible to measure the true primordial pattern in the CMB, because the dipole components $a_{1m}$ have been removed from the maps to compensate for the local motion relative to the local cosmic rest frame, including our nonlinear orbits within the galaxy and the Local Group. These motions are not known to nearly enough precision to separate the primordial dipole (Peebles, 2022). Nevertheless, a small fraction of the subtracted dipole is part of the intrinsic large-angle primordial pattern on spherical causal diamond surfaces and contributes to correlation in the angular range of causal shadows. Thus, a null shadow symmetry can only become apparent when the intrinsic portion of the dipole is included. If it is a true symmetry of the cosmic horizon, then there must exist a dipole that can be added to the observed CMB temperature map that realizes the symmetry.
The total correlation (Eq. 4) is a sum of even and odd Legendre polynomials, which are respectively symmetric and antisymmetric about $\Theta=\pi/2$. To produce zero correlation over a range symmetric around $\Theta=\pi/2$, no combination of even functions can cancel any combination of odd ones, so if an angular correlation function vanishes over a range $[\pi/2-\Theta_{0},\pi/2+\Theta_{0}]$ symmetric about $\Theta=\pi/2$, the even contributions and the odd contributions to the angular correlation function must vanish independently over that range.
This property allows a direct, model- and dipole- independent test of causal symmetry, that uses only even-parity correlation. In a band of angles symmetric around $\Theta=\pi/2$ determined by the 3D causal shadow (Eq. 15), the sum of even terms must vanish on its own, independently of any dipole or model parameters. The causal shadow of even-parity correlation is thus
$C_{\rm even}(|\Theta-\pi/2|<\arcsin[1/4])=0,$ (19)
or approximately $C_{\rm even}(75.52^{\circ}<\Theta<104.48^{\circ})=0$.
Furthermore, if the true primordial angular correlation including $\ell=1$ vanishes over an arbitrary range $[\alpha,\beta]$, then the sum of the even and odd Legendre polynomials in Eq. (4), measured only with $\ell>1$, departs from zero by a function of known form, the dipole harmonic term
$\mathcal{D}(\Theta)=\frac{3}{4\pi}C_{1}\cos(\Theta),$ (20)
where $C_{1}\geq 0$. Thus, if there is a causal shadow over a larger angular range (Eq. 17 or 18), the sum of the even and odd Legendre polynomials for $\ell>1$ must vanish after addition of a dipole of unknown amplitude.
4.2 Comparison with standard predictions
4.2.1 Even-parity and total correlation comparisons
We perform two types of comparisons with data. First, we directly compare the maps with the zero correlation predicted from causal coherence in the 3D causal shadow, using only even-parity correlation. Then, we use model-independent comparisons of data to explore whether maps are also consistent with zero total correlation over a larger range of angles, where the unmeasured dipole must be accounted for.
According to the causal shadow hypothesis, the even-parity correlation should vanish in the range of angles where all gravitational contributions to both odd and even contributions vanish (Eq. 19). As verified quantitatively by the rank comparison described below, the maps are indeed much closer to zero over this range than almost all realizations in the standard picture. Prediction and measurement in this comparison are model- and parameter-free.
Outside this range, the odd and even components do not separately vanish. The unmeasured dipole must be included to reveal any null symmetry, since odd-parity harmonics must be included. An added cosine function (Eq. 20) reproduces the effect of restoring any unobserved intrinsic dipole. The amplitude of this function is not known, which must be accounted for in statistical comparisons.
4.2.2 Standard realizations
To generate standard-model realizations, we used the Code for Anisotropies in the Microwave Background (CAMB) (Lewis & Challinor, 2011) to calculate $C_{\ell}^{\text{\,SM}}$ with the following six cosmological parameters from the Planck collaboration (Planck Collaboration, 2020a): dark matter density $\Omega_{c}h^{2}=0.120$; baryon density $\Omega_{b}h^{2}=0.0224$; Hubble constant $H_{0}=67.3$; reionization optical depth $\tau=0.054$; neutrino mass $m_{\nu}=0.06$ eV; and spatial curvature $\Omega_{k}=0.001$. For each realization, we calculated the angular power spectrum using Eq. 3. Then, we determined $C(\Theta)$ by summing Eq. 4 up to the sharp cutoff at $\ell_{\text{max}}=30$.
Correlation functions of realizations and CMB maps are shown in Fig. (1). On this scale, realizations with the same parameters display considerable cosmic variance. The 100 realizations shown for $C_{\rm even}$ directly illustrate examples of what would be expected in the standard picture.
Standard realized correlation functions include only $\ell>1$ harmonics. For the comparisons of total correlation shown in the second panel of Fig. (1), which include odd-parity harmonics, each realization is modified with a function of the form in Eq. (20) to minimize its residuals from zero. For realizations, this term does not have any relation to an intrinsic physical dipole: it is a “mock dipole” added to estimate how frequently the sum of $\ell>1$ harmonics in Eq. (4) comes as close to the maps as zero correlation in the posited range of angular separation. Although the residual variance of these comparisons exceeds that of purely even parity correlation in the 3D causal shadow, it appears that the measured $C(\Theta)$ in the posited range (Eqs. 17 or 18) is still closer to zero than almost all standard realizations, even when they have a mock dipole added.
4.2.3 Residuals
The striking visual impression of a null symmetry in the measured correlation can be verified quantitatively by a rank comparison of residuals. For angular power spectrum $C_{\ell}$, define the even-parity angular correlation function $C_{\text{\rm even}}(\Theta)$ as
$C_{\text{\rm even}}(\Theta)=\frac{1}{4\pi}\sum_{\begin{subarray}{c}\ell\;\!=\;\!2,\;\!4,\;\!6,\;\!...\end{subarray}}^{\ell_{\text{max}}}(2\ell+1){C}_{\ell}P_{\ell}(\cos\Theta),$ (21)
where $\ell_{\text{max}}=30$. Let $\{\Theta_{j,[\alpha,\beta]}\}_{i=1}^{N}$ denote a uniformly spaced lattice of points in the range $[\alpha,\beta]$. Then, define the even-parity residual
[…]
$\approx\sum_{i=1}^{N}[C_{\text{\rm even}}(\Theta_{i,[\alpha,\beta]})]^{2}\cdot\left(\frac{\beta-\alpha}{N}\right)\!,.$ (23)
Similarly, for the total residual, we define
[…]
$\approx\sum_{i=1}^{N}[\tilde{C}_{\beta}(\Theta_{i,[\alpha,\beta]})]^{2}\cdot\left(\frac{\beta-\alpha}{N}\right)\!,$ (25)
where
$\tilde{C}_{\beta}=C+\mathcal{D}_{\text{best-fit}}$
and $\mathcal{D}_{\text{best-fit}}$ is the dipole contribution that minimizes the residual.
We then define integrated variation residuals over three ranges of angles:
$\Delta_{\text{\rm even}}$ $\equiv\Delta_{\text{\rm even},[\pi/2-\Theta_{D}/2,\pi/2+\Theta_{D}/2]},$ (26)
$\Delta_{\text{minimal}}$ $\equiv\Delta_{\text{best-fit},[\pi/2-\Theta_{D}/2,2\pi/3]},$ (27)
$\Delta_{\text{maximal}}$ $\equiv\Delta_{\text{best-fit},[\pi/2-\Theta_{D}/2,3\pi/4]},$ (28)
where $\Theta_{D}/2=\arcsin(1/4)$ as derived above (Eq. 12). We use these three residuals as a measure of how compatible a given power spectrum $\left\{C_{\ell}\right\}_{\ell>1}$ is with the causal shadow symmetry. Each of the integrals must vanish for a power spectrum that exactly agrees with the causal shadow symmetry in the specified range. In practice, we found that $N=2000$ is a sufficiently high lattice resolution to approximate the integrals among different data sets and standard model realizations with negligible error.
Figure 4: Cumulative probability of deviations from zero correlation in standard realizations, compared with deviations of CMB maps, over various ranges of angular separation. The top panel directly compares deviations of even-parity correlation (Eq. 23) over the computed 3D causal shadow (Eq. 19). No parameters are used for this comparison. The middle panel compares deviations of total correlation (Eq. 25) over the minimal asymmetric shadow (Eq. 17), and the bottom panel compares deviations of total correlation over the maximal asymmetric shadow (Eq. 18); both of these allow for a mock-dipole correction to minimize residuals for each realization. In spite of variation between the maps, their variations are all much closer to zero than almost all standard realizations. The departures from zero are comparable with the differences between the maps, as expected if they are dominated by systematic measurement errors. In all of these comparisons, the Planck maps match zero better than the WMAP maps. In the cleanest direct comparison, which is represented by the top panel, the residuals $\Delta_{\text{\rm even}}$ found in the Planck maps are only 1 to 9 $\mu{\rm K}^{4}$, compared with $\Delta_{\text{\rm even}}\simeq 10^{4}\mu{\rm K}^{4}$ found in typical realizations. The probabilities for the standard picture to match such small values range from $10^{-4.3}$ to $10^{-2.8}$. The extraordinary overall precision of CMB isotropy can be characterized by the fractional measured dimensionless residual, which reaches its smallest value $\int C^{2}_{\rm even}/(2.7{\rm K})^{4}\sim 2\times 10^{-26}$ in the NILC map.
4.2.4 Rank comparison with standard realizations
Our three comparisons are as follows. First, we generate $N=2\cdot 10^{6}$ standard model realizations. We then evaluate $\Delta_{\text{\rm even}}(C(\Theta))$, $\Delta_{\text{minimal}}(C(\Theta))$, and $\Delta_{\text{maximal}}(C(\Theta))$ for these standard model realizations, such as those shown in Fig. (1), as well as the different measured CMB maps. For a given residual $\Delta_{\text{\rm even}}$, $\Delta_{\text{minimal}}$, or $\Delta_{\text{maximal}}$, the variation of this residual for different measured CMB maps gives a measure of the sensitivity of the residual to Galactic model uncertainties.
The top panel of Fig. (4) shows cumulative probability, the fraction of standard realizations with $\Delta_{\text{\rm even}}$ smaller than the value shown on the horizontal axis, and the vertical lines show the values of $\Delta_{\text{\rm even}}$ for different measured CMB maps. These numbers confirm the visual impression from Fig. (1): in the three Planck maps the value of $\Delta_{\text{\rm even}}$ ranges from 1 to 9 $\mu{\rm K}^{4}$, compared with values $\Delta_{\text{\rm even}}\simeq 10^{4}\mu{\rm K}^{4}$ found in typical realizations. Only a small fraction of standard realizations come as close to zero $\Delta_{\text{\rm even}}$ as the CMB temperature maps; the fraction in the Planck maps ranges from $10^{-2.8}$ to $10^{-4.3}$.
The middle panel shows the same quantities evaluated for $\Delta_{\text{minimal}}$, with the dipole-term adjustment described above. We again find that a small fraction of standard model realizations, ranging from about $10^{-1.8}$ (for WMAP) to $10^{-2.8}$ (for Commander), come as close to zero $\Delta_{\text{minimal}}$ as the measured CMB temperature maps. The values of $C_{1}$ for the best-fit dipole for the maps NILC, SMICA, COMMANDER, WMAP without its monopole, and WMAP with its monopole are approximately $365$, $341$, $322$, $392$, and $426$ $\mu$K2, respectively.
The lower panel shows the same quantities evaluated for $\Delta_{\text{maximal}}$, with the dipole-term adjustment described above. We again find that a small fraction of standard model realizations, ranging from about $10^{-2.3}$ (for WMAP) to $10^{-3.2}$ (for Commander), come as close to zero $\Delta_{\text{maximal}}$ as the measured CMB temperature maps. The values of $C_{1}$ for the best-fit dipole for the maps NILC, SMICA, COMMANDER, WMAP without its monopole, and WMAP with its monopole are approximately $384$, $390$, $389$, $440$, and $471$ $\mu$K2, respectively.
The total residuals $\Delta_{\text{minimal}}$ and $\Delta_{\text{maximal}}$ are significantly larger than the even residual $\Delta_{\text{\rm even}}$ over the narrower range of the 3D causal shadow. However, the variation between the maps is also larger, again consistent with the interpretation that departures from zero are due to inaccurate models of the Galaxy, and that intrinsic CMB correlations actually vanish.
To evaluate the sensitivity of this comparison to the chosen cutoff $\ell_{\text{max}}$, we repeated them for every $\ell_{\text{max}}$ value ranging from $25$ to $35$. For each value, the significance of our results, i.e. the fraction of standard model realizations having a residual as low as that of measured CMB temperature maps, changed only slightly, less by at least an order of magnitude than the variations in significance between different maps.
5 Interpretation
In standard inflation, correlation varies randomly among different realizations, and the tiny measured correlation must be interpreted as a statistical anomaly. Our rank comparison (Fig. 4) shows that the sky agrees with zero better than almost all standard realizations. The most direct comparison, as well as the smallest measured values of correlation, appear in even-parity correlation, where Planck maps show deviations from zero that are three to four orders of magnitude smaller than typical standard realizations. The standard interpretation is that this small correlation happens by chance, which is estimated to occur in a fraction $10^{-2.8}$ to $10^{-4.3}$ of realizations, depending on the map.
Another interpretation is that the nearly-zero correlation is due to initial conditions more symmetric than generally assumed. One possibility is an exact fundamental causal symmetry of quantum gravity that is not included in the standard model. The measured angular range of minimal correlation agrees with that derived here from a basic physical principle of causal coherence— that quantum fluctuations generate physical correlations only within causal diamonds bounded by inflationary horizons. In this interpretation, the small measured departure from zero correlation is attributed to measurement error, dominated by contamination by the Galaxy. This view is consistent with measured variation among the different maps.
The measurements appear to be consistent with the interpretation that CMB correlation on large angular scales preserves a signature of a new causal symmetry of initial conditions. The symmetry is not consistent with the standard effective quantum field theory approximation for formation of inflationary perturbations, but could arise from a deeper structure of quantum geometry that becomes significant on the scale of causal horizons.
6 Conclusion
The surprisingly small absolute value of the large-angle CMB correlation function has been known since the first measurements with COBE (Bennett et al., 1994; Hinshaw et al., 1996). Subsequent measurements from WMAP (Bennett et al., 2003, 2011) and then Planck (Planck Collaboration, 2020c) show values successively closer to zero. This anomaly and others have been thoroughly analyzed (e.g., (Copi et al., 2009; Muir et al., 2018)), but are not generally thought to present a compelling challenge to standard cosmological theory, both because the standard theory occasionally produces such small correlations by chance, and because there has not been a precisely formulated and physically compelling alternative expectation.
The significant new fact reported in this paper is that when the removed dipole component is accounted for, the magnitude of correlations is much smaller than previously documented. In particular, direct measurements of even-parity correlation in the Planck data at angular separations in the geometrically-calculated causal shadow are consistent with zero, with measured dimensionless fractional variations smaller than $10^{-25}$. Correlations are several orders of magnitude smaller than expected in standard cosmology, even though it starts with a perfectly isotropic unperturbed classical background; depending on the map, the probability of such small correlations occurring by chance in the Planck maps is $10^{-2.8}$ to $10^{-4.3}$.
We are thus led to suspect that nearly-zero large angle correlation may not be an extraordinary anomaly of our particular sky, but a signature of new physical symmetry in cosmic initial conditions, which are shaped by quantum gravity. As one example, we proposed a geometrically-defined symmetry that invokes causal, nonlocal “spooky” quantum entanglement of gravitational vacuum fluctuation states, which is not included in the standard theory of inflationary perturbations based on quantum field theory, but is a plausible feature for a deeper theory of quantum gravity that addresses entanglement with causal horizons. As discussed in the Appendix, a symmetry of this kind is broadly consistent with tests of classical concordance cosmology, which mainly (if not entirely) depend only on a nearly scale-invariant initial 3D power spectrum of perturbations averaged over large volumes. It is also consistent with all other experimental tests of QFT, none of which depend on quantized gravity.
If the causal-symmetry hypothesis is not true, it can be falsified by more precise measurement of nonzero correlations on the cosmic horizon within predicted causal shadows. The precision of the results reported here, and the significance of our null tests, are not limited by any fundamental source of noise, but by the accuracy of models of Galactic foreground emission. Tests of the symmetry can be improved with all-sky models of emission from the Galaxy that allow better measurements of the true CMB pattern on the largest scales. In this analysis, we have used the published all-sky, unmasked, Galaxy-subtracted CMB maps prepared by the satellite teams. We have not undertaken a close comparative study of the Galaxy models used to create those maps, but it is plausible that existing data from WMAP and Planck could be combined to create maps better optimized for studies of large-angle correlation symmetry. Galaxy models would improve with better spectral measurements, as proposed in new satellite concepts (Chluba et al., 2021). Other unique signatures of causal coherence are addressed briefly in the Appendix.
Data Availability
For data access from both WMAP and Planck (Planck Team, 2013, 2020), we acknowledge use of the Legacy Archive for Microwave Background Data Analysis (LAMBDA), part of the High Energy Astrophysics Science Archive Center (HEASARC), and the NASA/ IPAC Infrared Science Archive, which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration.
Acknowledgements
O.K. thanks the Cardiff Gravity Exploration Institute for its hospitality and acknowledges support from the UKRI Science and Technology Facilities Council (STFC) under grant ST/Y005082/1. We are also grateful for support from an anonymous donor.